Find the weight of H2 stored in 100-cm3 cylindrical tank at 100




oC and 100 atm (10, 132.5 KPa)




using (a) Ideal gas equation (b) van der Waals equation if a = 0.244 atm-L




2




/mole




2 and b = 0.0266




L/mole. (Note: Use the method of successive approximation to solve for the mass of H2




)

1
Expert's answer
2022-02-28T11:05:04-0500

V= 100cm3= 0.1L

T= 373k

P= 100atm

R= 10÷132.5= 0.0755

n= ?

n= PV/RT

n= (0.1×100)÷ (0.0755×373)

= 0.355